# Invariance of deep image quality metrics to affine transformations

### v1

@@ II-C Metric thresholds in physical units
Following the methodology in Figure 1, we can obtain the metric thresholds in physical units, $\theta_{\tau}^{M}$. For example, the rotation threshold is expressed in degrees. Through the function $g_{M}(\theta)$, specific to each metric, a value in physical units is assigned to the threshold in the common internal representation, computed as $\theta_{\tau}^{M}=g_{M}^{-1}(\mathcal{D}_{\tau})$. These values can be compared numerically with the thresholds obtained for humans, as they are expressed in the same units because they were obtained with the psychometric functions, Section 3. Therefore, a metric can be said to have human behavior if the invisibility threshold for a certain affine transformation coincides, $\theta_{\tau}^{H}==\theta_{\tau}^{M}$. In particular, since $\mathcal{D}_{\tau}$ has an associated uncertainty, $\theta_{\tau}^{M}$ also has a confidence interval. In this case, we will check whether the human threshold falls within the confidence interval of each metric.
@@ II-D Sensitivities of metrics and humans for the different distortions
The sensitivity for a small distortion is usually defined as the inverse of the energy required to be above the invisibility threshold (Eq. 5), i.e, expressing the distortion $\theta_{\tau}^{H}$ in RMSE units. In the case of metrics, this general definition reduces to the derivative of the transduction function concerning the energy of the distortion, i.e. its slope.
@@
In this way, we can obtain the sensitivity of a metric for each affine transform, order them and check if the order matches the human ones. A real example of this can be found in Figure 7 (the rest of the figures can be found in Appendix LABEL:app:apndD), where the vertical lines represent the human thresholds expressed in terms of energy and the different curves describe the behavior of that metric for the various transformations. Given two curves, the most sensitive will be the one with a higher slope.
@@ III-A Affine Transformations
An affine transformation or affine application (also called affinity) between two affine spaces is a transformation that satisfies Equation 6.
@@
Some examples of affine transformations are geometric contraction, expansion, dilation, reflection, rotation, or shear. In this work, we are going to focus specifically on translation, rotation scale, and illuminant changes. For each tested affine transformation, the original images are modified in the following ways:
@@
  • Illuminant changes: We have desaturated the original images and modified the illuminant for 20 hues in 8 saturations. Following the distribution at the locus as indicated in Figure 8.
@@ IV Results
TABLE III: Results associated with the least demanding test, on the order of the sensitivities and their comparison with the human order. The first row of the table shows the order of the humans, which does not change depending on the database. Displayed in two different colors: In bold, the order of the geometric transformations and, in italic, the order of the YB and RG illuminants.
@@ V Discussion
Getting into Table III, it is not surprising to see that the ordering of the geometric transformations does not depend on the dataset chosen for the human observers. Regarding the metrics’ performance, PerceptNet is the only metric that almost maintains the ordering of being more sensitive to geometric over chromatic transformations. All the metrics are generally more sensitive to YB than RG, which matches human performance. All in all, no metric can be said to have human-like behavior even with the least strict test.

### v2

@@ II-C Metric thresholds in physical units
Following the methodology in Figure 1, we can obtain the metric thresholds in physical units, $\theta_{\tau}^{M}$. For example, the rotation threshold is expressed in degrees. Through the function $g_{M}(\theta)$, specific to each metric, a value in physical units is assigned to the threshold in the common internal representation, computed as $\theta_{\tau}^{M}=g_{M}^{-1}(\mathcal{D}_{\tau})$. These values can be compared numerically with the thresholds obtained for humans, as they are expressed in the same units because they were obtained with the psychometric functions, Section 3. Therefore, a metric can be said to have human behavior if the invisibility threshold for a certain affine transformation coincides, $\theta_{\tau}^{H}=\theta_{\tau}^{M}$. In particular, since $\mathcal{D}_{\tau}$ has an associated uncertainty, $\theta_{\tau}^{M}$ also has a confidence interval. In this case, we will check whether the human threshold falls within the confidence interval of each metric.
@@ II-D Sensitivities of metrics and humans for the different distortions
The sensitivity for a small distortion is usually defined as the inverse of the energy required to be above the invisibility threshold, i.e, expressing the distortion $\theta_{\tau}^{H}$ in RMSE units. In the case of metrics, this general definition reduces to the derivative of the transduction function concerning the energy of the distortion, i.e. its slope [28].
@@
In this way, we can obtain the sensitivity of a metric for each affine transform, order them and check if the order matches the human ones. A real example of this can be found in Figure 7 (the rest of the figures can be found in Appendix D), where the vertical lines represent the human thresholds expressed in terms of energy and the different curves describe the behavior of that metric for the various transformations, where both are calculated as the mean over the whole database. Given two curves, the most sensitive will be the one with a higher slope.
@@ III-A Affine Transformations
An affine transformation or affine application (also called affinity) between two affine spaces is a transformation that satisfies:
@@
Some examples of affine transformations that follow Equation 7 are geometric contraction, expansion, dilation, reflection, rotation, or shear; and some examples following Equation 8 are changes in contrast, luminance, and illuminant. In this work, we are going to focus specifically on translation, rotation scale, and illuminant changes. For each tested affine transformation, the original images are modified in the following ways:
@@
  • Illuminant changes: We have desaturated the original images and modified the illuminant for 20 hues in 8 saturations, i.e. angle with respect to the $x$ axis, and distance with respect to the central white point respectively. Following the distribution at the locus as indicated in Figure 8.
@@ IV Results
TABLE III: Results associated with the least demanding test, on the order of the sensitivities and their comparison with the human order. The first row of the table shows the order of the humans, which does not change depending on the database. Displayed in two different colors: In purple, the order of the geometric transformations and, in pink, the order of the YB and RG illuminants.
@@ V Discussion
Getting into the metrics’ performance regarding the ordering of the transformations from Table III, PerceptNet is the only metric that almost maintains the ordering of being more sensitive to geometric over chromatic transformations. All the metrics are generally more sensitive to YB than RG, which matches human performance in energy definition. All in all, no metric can be said to have human-like behavior even with the least strict test.
@@ Acknowledgment
Appendix A.
The following figures show examples from the TID13 database sorted by DMOS for which the reader can make a first estimate of the threshold value of human invisibility for the DMOS, i.e., a first estimation of $\mathcal{D}_{\tau}$. For each reference image (at the top) some distorted images are selected, sorted in increasing order by the DMOS value. With these images, it can be tested, in a rather coarse way, which DMOS value tells us what is invisible and what is not to the observer. Roughly, the value is around 0.45, since the second image is usually indistinguishable from the original but the third one always differs.
Appendix B.
Figure 13 shows the results for the all databases for the geometric affine transformations (translation, rotation and scale). In addition, the first subgraph shows again the proposed methodology with the cuts on the y-axis, marked in gold color; it is the only one that also shows the legend and it is shared by all the others.
Figure 13: Results of the different metrics in the experiments for the geometric affine transformations. On these graphs we will apply what we have seen in Figure 1 to create intervals to compare with the human thresholds. In the different columns, we have the results differentiated by database (and in the rows, by affine transformation (translation, rotation and scale).
Appendix C.
Figures 16, 17, 15, 14 show the results for the all databases for the illuminant transformations showing the final ellipse in the spectral locus. The title of each subfigure specifies the errors made with respect to the MacAdam ellipse at that point in space (left) and with respect to the experimentally obtained ellipse (right).
Figure 14: Results of the different metrics in the color discrimination experiments, together with their error for ImageNet data. The error with respect to the fitted MacAdam ellipse is shown on the left whereas the error with respect to the experimentally obtained ellipse is shown on the right. The first subfigure also shows the direction of the 20 hues. The metric that makes the least error with respect to the MacAdam ellipses is PIM in both cases.
Figure 15: Results of the different metrics in the color discrimination experiments, together with their error for TID13 data. The error with respect to the fitted MacAdam ellipse is shown on the left whereas the error with respect to the experimentally obtained ellipse is shown on the right. The first subfigure also shows the direction of the 20 hues. The metric that makes the least error with respect to the MacAdam ellipses is PIM in both cases.
Figure 16: Results of the different metrics in the color discrimination experiments, together with their error for MNIST data. The error concerning the fitted MacAdam ellipse is shown on the left whereas the error with respect to the experimentally obtained ellipse is shown on the right. The first subfigure also shows the direction of the 20 hues. The metric that makes the least error with respect to the MacAdam ellipses is DISTS in both cases.
Figure 17: Results of the different metrics in the color discrimination experiments, together with their error for ImageNet data. The error with respect to the fitted MacAdam ellipse is shown on the left whereas the error with respect to the experimentally obtained ellipse is shown on the right. The first subfigure also shows the direction of the 20 hues. The metric that makes the least error with respect to the MacAdam ellipses is DISTS in both cases.
Appendix D.
Figures 18, 19,20, and 21 show the results for the all databases for the comparison of sensibilities. The vertical lines show the human thresholds for each transformation, the different curves, the behaviour of each model. The order of the vertical lines is compared with the value of the slope. A steep slope implies higher sensitivity.
Figure 18: Sensitivity test results for the MNIST dataset. Table III shows in detail the order of the sensitivities together with the human ordering, where it can be seen that no metric follows the order correctly. However, in all cases the relative order of illuminant changes is respected, i.e. they are more sensitive to the RG direction than to the YB direction.
Figure 19: Sensitivity test results for the CIFAR10 dataset. Table III shows in detail the order of the sensitivities together with the human ordering, where it can be seen that no metric follows the order correctly.
Figure 20: Sensitivity test results for the TID13 dataset. Table III shows in detail the order of the sensitivities together with the human ordering, where it can be seen that no metric follows the order correctly.
Figure 21: Sensitivity test results for the ImageNet dataset. Table III shows in detail the order of the sensitivities together with the human ordering, where it can be seen that no metric follows the order correctly.